OpenStudy (anonymous):

Solve. 8 – 3x – x = 12 x= ?

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

-3x-x=?

OpenStudy (anonymous):

make it so that you are adding a negative so it would look like this 8+-3x-x=12

OpenStudy (anonymous):

-4x = 12-8 -4x = 4 x = -4/4 x=-1

OpenStudy (anonymous):

Thanks :)